# Fast Gradient Computation for RoPE Attention in Almost Linear Time

### v2

@@ 3.2 Problem Definition
3.3 Polynomial Approximation of Exponential
@@
3.4 Time Complexity of Multiplications
@@
3.5 $\mathsf{SETH}$ Hypothesis
@@
Here, we introduce the Strong Exponential Time Hypothesis ($\mathsf{SETH}$):
Hypothesis 3.7 ($\mathsf{SETH}$).
@@ 3.6 Basic Facts
Fact 3.8 (Folklore).
@@
Fact 3.9 (Inner Products).
@@
Fact 3.10 (Tensor trick).
Fact 3.11.
  • $(\exp(A_{1}XA_{2}^{\top})_{j_{0},*})^{\top}=\exp(\mathsf{A}_{j_{0}}x)$
  • $\operatorname{vec}(\exp(A_{1}XA_{2}^{\top}))=\exp(\mathsf{A}x)$,
Proof.
From Lemma and Def. 3.10, we are able to prove this fact. ∎
3.7 Useful Definitions
@@
Definition 3.13 (Diagonal matrix $\alpha(x)$).
@@
Then, we define $s(x)$ for the normalized softmax $(D^{-1}\cdot\mathrm{softmax})$.
Definition 3.14 (Normalized softmax $s(x)$).
From Def. 3.12, it defines $u(\cdot)_{j_{0}}$ , and we have $\alpha(\cdot)_{j_{0}}$based on Def. 3.13. The function $s(x)_{j_{0}}$ maps a $d^{4}$-dimensional vector to an $n$-dimensional vector given every $j_{0}\in[n]$ such that
$s(x)_{j_{0}}:=\alpha{(x)}_{j_{0}}^{-1}u(x)_{j_{0}}.$
@@
Definition 3.15 (Value matrix $v(y)$).
Let $A_{3}\in\mathbb{R}^{n\times d}$ be a matrix. We define $v(y)_{i_{0}}$ as the $i_{0}$-th column of $v(y)$. We define the function $v(y)_{i_{0}}$ maps a $d^{2}$-dimensional vector to an $n$-dimensional vector, given each $i_{0}$ in the set $[d]$, such that
$v(y)_{i_{0}}:=A_{3}Y_{*,i_{0}}$
where $y\in\mathbb{R}^{d^{2}}$ is the vectorization of $n\times n$ matrix $Y$.
@@
Definition 3.16.
From Def. 3.14, with every $j_{0}$ in the set $[n]$, it gives $s(x)_{j_{0}}$ as an $n$-dimensional normalized vector, and we define $v(y)_{i_{0}}$ based on Def. 3.15 given that each $i_{0}\in[d]$. Defining a function $\ell(x)_{j_{0},i_{0}}$ maps a $d^{4}$-dimensional vector to a scalar with each $j_{0}\in[n]$ and each $i_{0}\in[d]$ such that
@@
Definition 3.17.
@@
3.8 Reformulation of the Loss Function
In this section, we are able to reformulate and simplify the loss function based on the definitions provided in Section 3.7. This reformulation provides a structured representation of the loss in terms of its components, using the tensor trick to simplify computations and facilitate analysis.
Lemma 3.18.
With the following conditions
  • Given three $n\times d$ input sequence matrices $A_{1}$, $A_{2}$, and $A_{3}$,
  • Let $\mathsf{A}=A_{1}\otimes A_{2}\in\mathbb{R}^{n^{2}\times d^{2}}$ and $\mathsf{X}=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$, where $\otimes$ denotes Kronecker product.
  • Given $W$ is a $n^{2}\times d^{2}$ matrix, we define $\widetilde{A}=A\oslash W$. Let $j_{0}\in[n]$, and we define $\widetilde{A}_{j_{0}}\in\mathbb{R}^{n\times d^{2}}$ be a block of size $n\times d^{2}$ from $\widetilde{A}$.
  • $E\in\mathbb{R}^{n\times d}$ be a matrix. For $j_{0}\in[n]$ and $i_{0}\in[d]$, we define $E_{j_{0},i_{0}}$ as the $(j_{0},i_{0})$-th position or entry the $n\times d$ matrix $E$.
  • Here are three squared matrices $X_{1},X_{2}\in\mathbb{R}^{d\times d}$, let $Y\in\mathbb{R}^{d\times d}$
  • Based on Def. 3.2, we have $\mathsf{Loss}(\mathsf{X})$.
  • Based on Def. 3.17, For $j_{0}$ in the set $[n]$ and $i_{0}$ in the set[d]$[d]$, we get $\mathsf{Loss}(\mathsf{X})_{j_{0},i_{0}}$
Then, we have
Proof.
$=$ $~{}\sum_{j_{0}=1}^{n}\sum_{i_{0}=1}^{d}0.5\cdot(\langle\langle\exp(\widetilde{\mathsf{A}}_{j_{0}}x),\mathbf{1}_{n}\rangle^{-1}\exp(\widetilde{\mathsf{A}}_{j_{0}}x),A_{3}Y_{*,i_{0}}\rangle-E_{j_{0},i_{0}})^{2}$
$=$ $~{}\sum_{j_{0}=1}^{n}\sum_{i_{0}=1}^{d}\mathsf{Loss}(x)_{j_{0},i_{0}}$
where the 1st equality is based on Def. 3.2, the definition of Frobenius norm derives the 2nd equality, the 3rd equality is due to Def. 3.14 and Def. 3.15, and the 4th step is based on Def. 3.17. ∎
4 RoPE Attention Gradient Calculation

### v3

@@ 3.2 Problem Definition
3.3 Reformulation of the Loss Function
In this section, we are able to reformulate and simplify the loss function based on the definitions provided in Section B. This reformulation provides a structured representation of the loss in terms of its components, using the tensor trick to simplify computations and facilitate analysis.
Lemma 3.4 (Loss Function Formulation).
[…]
Proof.
$=$ $~\sum_{j_{0}=1}^{n}\sum_{i_{0}=1}^{d}0.5\cdot(\langle\langle\exp(\widetilde{\mathsf{A}}_{j_{0}}x),\mathbf{1}_{n}\rangle^{-1}$
$~\exp(\widetilde{\mathsf{A}}_{j_{0}}x),A_{3}Y_{*,i_{0}}\rangle-E_{j_{0},i_{0}})^{2}$
$=$ $~\sum_{j_{0}=1}^{n}\sum_{i_{0}=1}^{d}\mathsf{Loss}(x)_{j_{0},i_{0}}$
where the 1st equality is based on Def. 3.2, the definition of Frobenius norm derives the 2nd equality, the 3rd equality is due to Def. B.3 and Def. B.4, and the 4th step is based on Def. B.6. ∎
4 Exact Gradient Computation Time
In this section, we provide the gradient computations of RoPE attentions. In Section 4.1, we formulate the gradient in its closed form. In Section 4.2, we conduct a time complexity analysis on the exact computation of RoPE attention gradients.
4.1 Reformulate the Gradient into Its Closed Form
In this section, we present the closed-form gradient of RoPE attention.
Lemma 4.1 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}$, Informal Version of Lemma C.4).
For every $i\in[d^{4}]$, we choose the following functions
  • The Normalized Softmax function $s(x)_{j_{0}}\in\mathbb{R}^{n}$ (see Definition B.3),
  • The Error term $\ell(x)_{j_{0},i_{0}}\in\mathbb{R}$ (see Definition B.5),
  • The Loss term $\mathsf{Loss}(x)_{j_{0},i_{0}}\in\mathbb{R}$ (see Definition B.6),
  • We define $\beta(x)_{j_{0}}\in\mathbb{R}^{n}$ is $\underbrace{A_{3}Y}_{n\times d}\underbrace{\ell(x)_{j_{0},*}^{\top}}_{d\times 1}$
  • We define $\gamma(x)_{j_{0}}\in\mathbb{R}^{n}$ is $(\operatorname{diag}(s(x)_{j_{0}})-s(x)_{j_{0}}s(x)_{j_{0}}^{\top})\beta(x)_{j_{0}}$
Then, we get $\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}=\underbrace{\widetilde{\mathsf{A}}^{\top}}_{d^{4}\times n^{2}}\operatorname{vec}(\underbrace{\gamma(x)}_{n\times n}).$
Proof.
See full proof at Lemma C.4. ∎
4.2 Time Complexity for Computing the Gradient of RoPE Attention
In this section, we provide the time complexity of computing the exact gradient of RoPE attention.
Theorem 4.2 (RoPE attention gradient computation time complexity).
We define three $n\times d$ input matrices as $A_{1},A_{2},A_{3}$, and the $n\times d$ approximated attention computation matrix as $E$. We define several input fixed matrices as $X_{1},X_{2},Y\in\mathbb{R}^{d\times d}$. We define $\mathsf{X}=X_{1}\otimes X_{2}$, $\mathsf{A}=A_{1}\otimes A_{2}$. We define $x:=\operatorname{vec}(\mathsf{X})$ and try to get the Loss function gradient. Let $g:=\frac{\mathrm{d}\mathsf{Loss}(X_{1},X_{2})}{\mathrm{d}x}$ where $\mathsf{Loss}(X_{1},X_{2})$ from Def. 3.3. Then, it costs $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ time to get the gradient $g\in\mathbb{R}^{d^{4}}$.
Proof.
See full proof at Theorem C.8. ∎
Note that $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))\geq\Omega(n^{2})$. Thus, the naive RoPE attention gradient computation is a complexity obstacle in practice, as discussed in Section 1. Based on the closed formulation in Lemma 4.1, we derive our acceleration method, which will be introduced in the following section.
5 Compute RoPE Attention Gradient in Almost Linear Time
In Section 5.1, we discuss the techniques we used to develop the almost linear time algorithm. In Section 5.2, we provide the proof of approximating $s(x)$ in almost linear time. In Section 5.3, we give the proof to approximate the error term $\ell(x)$. In Section 5.4, we show how to approximate $\beta(x)$ in almost linear time. In Section 5.5, we present our technique to approximate $\gamma(x)$. In Section 5.6, we present our main results, which compute RoPE Attention gradient in almost linear time.
5.1 Technique Overview
Lemma 5.1 (Theorem 1.3 in [5]).
Recall that the closed form gradient of RoPE attention is $\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}=\widetilde{\mathsf{A}}^{\top}\operatorname{vec}(\gamma(x))$ from Lemma 4.1. We need to show $\gamma(x)$ can be low-rank approximated in $O(n^{1+o(1)})$ time with $1/\operatorname{poly}(n)$ error.
To low rank approximate $\gamma(x)$, we use the strategy to split $\gamma(x)$ into two terms, $\gamma_{1}(x)$ and $\gamma_{2}(x)$, and run the approximation separately. From Lemma 4.1, $\gamma(x)_{j_{0}}\in\mathbb{R}^{n}$ is $(\operatorname{diag}(s(x)_{j_{0}})-s(x)_{j_{0}}s(x)_{j_{0}}^{\top})\beta(x)_{j_{0}}$. We define $\gamma_{1}(x)_{j_{0}}=\operatorname{diag}(s(x)_{j_{0}})\beta(x)_{j_{0}}$and $\gamma_{2}(x)_{j_{0}}=s(x)_{j_{0}}s(x)_{j_{0}}^{\top}\beta(x)_{j_{0}}$; thus, we can have $\gamma(x)=\gamma_{1}(x)-\gamma_{2}(x)$.
In the definitions of $\gamma_{1}(x)$ and $\gamma_{2}(x)$ provided above, they both contain $s(x)$ and $\beta(x)$. In order to find the almost linear time algorithm of $\gamma(x)$, we need to first show that there exists $O(n^{1+o(1)})$ time complexity approximation for $s(x)$ and $\beta(x)$ with $\epsilon/\operatorname{poly}(n)$ error first. From Lemma 4.1, we have $\beta(x)_{j_{0}}\in\mathbb{R}^{n}$ is $A_{3}Y\ell(x)_{j_{0},*}^{\top}$. Based on the $\beta(x)$ definition, we need to show $\ell(x)$ can be approximated in almost linear time first.
Overall, to develop the $O(n^{1+o(1)})$ time complexity algorithm to compute RoPE gradients with $\epsilon/\operatorname{poly}(n)$ error, we need to prove the existence of almost linear time algorithms for $s(x),\ell(x),\gamma(x)$, and $\beta(x)$ with low rank approximation.
5.2 Approximate $s$ Using Low Rank Approximation
In this section, we use the low-rank approximation technique to approximate the normalized Softmax $s(x)$ in almost linear time.
Lemma 5.2 (Low Rank Approximate $s(x)$).
Proof.
By definition of $A(\mathsf{X})$, we have $\operatorname{vec}(A(\mathsf{X}))=\exp(\mathsf{A}\oslash\mathsf{W})\operatorname{vec}(X).$
Hence, using the tensor trick, we have
We can reformulate the attention matrix $A$ as, for $i,j\in[n]$
Thus, we can show that $\mathsf{A}_{i+(j-1)n}\mathsf{X}w_{i-j}=(A_{1,i,*}\otimes A_{2,j,*})(X_{1}\otimes X_{2})\operatorname{vec}(W_{i-j})=A_{1,i,*}X_{1}W_{i-j}X_{2}^{\top}A^{\top}_{2,j,*},$ where 1st equality uses definitions of $w_{i-j},\mathsf{A}$, and $\mathsf{X}$, and the second step uses the tensor trick. We complete our proof after applying Lemma 5.1. ∎
5.3 Approximate $\ell$ Using Low Rank Approximation
Lemma 5.3 (Low Rank Approximate $\ell(x)$).
Let $d$ equal $O(\log n)$. Suppose we can use $O(\log n)$ bits to write every entry in $E,v(y)\in\mathbb{R}^{n\times d}$. Define the $\ell(x)\in\mathbb{R}^{n\times d}$ as specified in Def. B.5. Then, we have $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$ such that $\|U_{1}V_{1}^{\top}v(y)-E-\ell(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n).$
Proof.
Here, we present the bound as follows.
$=$ $~\|v(y)\|_{\infty}\cdot\|U_{1}V_{1}^{\top}-s(x)\|_{\infty}$
$\leq$ $~\epsilon/\operatorname{poly}(n),$
where the 1st is because of Def. B.5, 2nd step is based on the distributive law, and 3rd step is due to Lemma 5.2. ∎
5.4 Approximate $\beta$ Using Low Rank Approximation
Lemma 5.4 (Low Rank Approximate $\beta(x)$).
Let $k_{2}=n^{o(1)}$. We define $\ell(x)\in\mathbb{R}^{n\times d}$ based on Def. B.5, and $v(y)\in\mathbb{R}^{n\times d}$ based on Def. B.4. We suppose $\beta(x)$ is equal to $v(y)\ell(x)^{\top}$, which is an $n\times n$ matrix. Let $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ such that $\|U_{2}V_{2}^{\top}-\beta(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$. In $n^{1+o(1)}$ time, we can get $U_{2},V_{2}$.
Proof Sketch.
We define $\widetilde{\beta}(x)\approx\beta(x)$ and $\widetilde{\beta}(x)=v(y)v(y)^{\top}V_{1}U_{1}^{\top}-v(y)E^{\top}$. We can first compute $v(y)^{\top}V_{1}$ as it can be computed in $n^{1+o(1)}$ time. Given all low rank matrices, we can have $U_{2},V_{2}$ where $k_{2}=\max\{d,k\}+d=n^{o(1)}$. Then we can compute $\|\widetilde{\beta}(x)-\beta(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$. (See full proof at Lemma D.3) ∎
5.5 Approximate $\gamma$ Using Low Rank Approximation
First, we show the low-rank approximation of $\gamma_{1}(x)$.
Lemma 5.5 (Low Rank Approximate $\gamma_{1}(x)$).
Proof Sketch.
Let $U_{3}=U_{1}\oslash U_{2}$ and $V_{3}=V_{1}\oslash V_{2}$, and we can use $n^{1+o(1)}$ time to get them. From Lemma 5.2 and Lemma 5.4, we have $\widetilde{s}(x)=U_{1}V_{1}^{\top}$ and $\widetilde{\beta}(x)=U_{2}V_{2}^{\top}$. Based on Fact A.5, we can compute $\|U_{3}V_{3}^{\top}-\gamma_{1}(x)\|_{\infty}\leq\|U_{3}V_{3}^{\top}-\operatorname{diag}(s(x))\beta(x)\|_{\infty}\leq\frac{\epsilon}{\operatorname{poly}(n)}$ (See full proof at Lemma D.4). ∎
Next, we show the low-rank approximation of $\gamma_{2}(x)$.
Lemma 5.6 (Low Rank Approximate $\gamma_{2}(x)$).
Proof Sketch.
Let $\rho(x)\in\mathbb{R}^{n}$ be defined by $\rho(x)_{j_{0}}=s(x)_{j_{0}}\beta(x)_{j_{0}}$. We construct $\widetilde{\rho}(x)$ so that $(U_{1}V_{1})_{j_{0},*}^{\top}\approx s(x)_{j_{0}}$ and $(U_{2}V_{2})_{j_{0},*}^{\top}\approx\beta(x)_{j_{0}}$, implying $\widetilde{\rho}(x)_{j_{0}}=(U_{1}V_{1})_{j_{0},*}\cdot(U_{2}V_{2})_{j_{0},*}^{\top}$. Pre-computing $V_{1}V_{2}^{\top}$ takes $n^{1+o(1)}$ time, and then computing each $\widetilde{\rho}(x)_{j_{0}}$ costs $O(k_{1}k_{2})$, giving a total of $O(nk_{1}k_{2})=n^{1+o(1)}$. Next, we approximate $s(x)$ by $\widetilde{s}(x)=U_{1}V_{1}^{\top}$ and define $\widetilde{\gamma}_{2}(x)=\widetilde{s}(x)\,\operatorname{diag}(\widetilde{\rho}(x))$; with $U_{4}=U_{1}$ and $V_{4}=\operatorname{diag}(\widetilde{\rho}(x))\,V_{1}$, we have $\widetilde{\gamma}_{2}(x)=U_{4}V_{4}^{\top}$. To bound the error, note that $\|\widetilde{\gamma}_{2}(x)-\gamma_{2}(x)\|_{\infty}=\max_{j_{0}}\|\widetilde{s}(x)_{j_{0}}\,\widetilde{\rho}(x)_{j_{0}}-s(x)_{j_{0}}\,\rho(x)_{j_{0}}\|_{\infty}$ can be split and bounded via the triangle inequality so that $\|\widetilde{s}(x)_{j_{0}}-s(x)_{j_{0}}\|_{\infty}$ and $\|\widetilde{\rho}(x)_{j_{0}}-\rho(x)_{j_{0}}\|_{\infty}$ are small, leading to an overall error of at most $\epsilon/\operatorname{poly}(n)$, which completes the proof. (See full proof at Lemma D.5) ∎
5.6 Fast Computation in Almost Linear Time
Based on Section 5.1, we have proved the almost linear time approximation of $s(x),\ell(x),\gamma(x)$, and $\beta(x)$ in Lemma D.1D.2D.3D.5, and D.4. We are now ready to show our main result, which is to approximate RoPE gradient computation in almost linear time.
Assuming the entries of $A_{1},A_{2},X_{1},X_{2},Y,E$ are represented using $O(\log n)$ bits, there is an $n^{1+o(1)}$ time algorithm to solve $\mathsf{AAttLGC}(n,d=O(\log n),B=o(\sqrt{\log n}))$, from Def. 3.3, with the accuracy upper bounded by $\frac{1}{\operatorname{poly}(n)}$ . To be more specific, a gradient vector $\widetilde{g}\in\mathbb{R}^{d^{4}}$ comes out of our algorithm where $\|\frac{\mathrm{d}\mathsf{Loss}}{\mathrm{d}x}-\widetilde{g}\|_{\infty}\leq\frac{1}{\operatorname{poly}(n)}$.
Proof Sketch.
By Lemma D.5 and Lemma D.4, there exist matrices $\gamma_{1}(x)$ and $\gamma_{2}(x)$ such that $\gamma(x)=\gamma_{1}(x)-\gamma_{2}(x)$. We assume these lemmas follow from the low-rank approximations in Lemmas D.1D.3, allowing us to write $\widetilde{\gamma}_{1}(x)=U_{3}V_{3}^{\top}$ and $\widetilde{\gamma}_{2}(x)=U_{4}V_{4}^{\top}$ in $n^{1+o(1)}$ time. From Lemma 4.1, the reformulated gradient is $\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}=\widetilde{\mathsf{A}}^{\top}\operatorname{vec}(\gamma(x))$, and hence the total running time remains $n^{1+o(1)}$. To bound the error, we show that
$\|\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}-\widetilde{g}\|_{\infty}$ $~=\|\widetilde{\mathsf{A}}^{\top}(\operatorname{vec}(\gamma(x))-\operatorname{vec}(\widetilde{\gamma}(x)))\|_{\infty}$
$~\leq\|\widetilde{\mathsf{A}}^{\top}\|_{\infty}\,\|\gamma(x)-\widetilde{\gamma}(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n),$
This completes the proof. (See full proof at Theorem D.6) ∎
6 Hardness
In this section, we provide the lower bound results to compute the gradient of RoPE attention. The hardness result shows that under the widely accepted $\mathsf{SETH}$, the bounded entries condition is necessary for achieving subquadratic runtime.
Theorem 6.1 (Lower bound, informal version of Theorem E.1).
Assuming $\mathsf{SETH}$, for any $q>0$, for the $\mathsf{ARAttLGC}(n,d=O(\log n),B=\omega(\sqrt{\log n})$, there does not exist an algorithm which can be executed in time $O(n^{2-q})$ based on Def. 3.3.
Proof.
See the full proof at Theorem E.1. ∎
In Theorem 6.1, we show that under the Strong Exponential Time Hypothesis (SETH) (see Hypothesis A.4), computing the gradient of RoPE attention remains computationally hard. Specifically, for any constant $q>0$, no algorithm can compute the gradient in time $O(n^{2-q})$ when $d=O(\log n)$ and $B=\omega(\sqrt{\log n})$. This result establishes a lower bound that fundamentally limits the efficiency of gradient computation for RoPE attention.
7 Conclusion
This paper presents the first efficient backward gradient computation, assuming bounded entries for the RoPE-based attention mechanism. We achieve almost linear time complexity by leveraging polynomial methods and the Fast Fourier Transform, making the forward and backward computations comparably efficient. Additionally, we demonstrate that conditions exist under which performance better than quadratic can be realized, consistent with the lower bounds suggested by the Strong Exponential Time Hypothesis (SETH).
These findings not only improve the computational efficiency of RoPE-based attention mechanisms but also provide a foundation for exploring sub-gradient computations in other advanced attention variants of neural networks. This work highlights the connection between algorithm design and computational complexity theory, unveiling new possibilities for the development of large transformer models. Future research could extend these results to cases involving unbounded entries and assess the real-world implications of these theoretical advancements for large language models. Furthermore, applying this approach to other positional encoding mechanisms could further enhance the scalability of state-of-the-art transformer models.
Appendix
Appendix A Preliminary
In Section A.1, we talk about polynomial approximation of the exponential function. In Section A.2, we talk about the time complexity of matrix multiplications, setting up the framework for analyzing efficiency in attention computation. In Section A.3, we talk about the Strong Exponential Time Hypothesis (SETH). In Section A.4, we talk about mathematical properties and tricks, such as the tensor trick and row-wise Kronecker products, which enable efficient matrix-vector operations.
A.1 Polynomial Approximation of Exponential
@@
A.2 Time Complexity of Multiplications
@@
A.3 $\mathsf{SETH}$ Hypothesis
@@
Our results are built on the common conjecture. [16] introduce the Strong Exponential Time Hypothesis (SETH) as a stronger form of the $\mathsf{P}\neq\mathsf{NP}$ conjecture. It suggests that our current best $\mathsf{SAT}$ algorithms are optimal and is a popular conjecture for proving fine-grained lower bounds for a wide variety of algorithmic problems [12, 27].
Hypothesis A.4 ($\mathsf{SETH}$).
@@ A.4 Basic Facts
Fact A.5 (Row-wise Kronecker product).
@@
Fact A.6 (Tensor trick).
Fact A.7.
  • $(\exp(A_{1}XA_{2}^{\top})_{j_{0},*})^{\top}=\exp(\mathsf{A}_{j_{0}}x)$
  • $\operatorname{vec}(\exp(A_{1}XA_{2}^{\top}))=\exp(\mathsf{A}x)$,
Proof.
From Lemma and Def. A.6, we are able to prove this fact. We omit the details here since the proof is straightforward. ∎
Fact A.8.
We suppose there are three vectors of $n$ dimension $x,y,z$. Thus, we get
  • $\langle x\circ y,z\rangle=x^{\top}\operatorname{diag}(y)z$.
  • $\langle x,y\rangle=\langle x\circ y,{\bf 1}_{n}\rangle$.
@@
Fact A.9 (Inner Products).
@@
Appendix B Key Definitions of RoPE Attention
@@
Definition B.2 (Diagonal matrix $\alpha(x)$).
@@
We define $s(x)$ for the normalized softmax $(D^{-1}\cdot\mathrm{softmax})$.
Definition B.3 (Normalized softmax $s(x)$).
From Def. B.1, it defines $u(\cdot)_{j_{0}}$ , and we have $\alpha(\cdot)_{j_{0}}$based on Def. B.2. The function $s(x)_{j_{0}}$ maps a $d^{4}$-dimensional vector to an $n$-dimensional vector given every $j_{0}\in[n]$ such that $s(x)_{j_{0}}:=\alpha{(x)}_{j_{0}}^{-1}u(x)_{j_{0}}.$
@@
Definition B.4 (Value matrix $v(y)$).
Let $A_{3}\in\mathbb{R}^{n\times d}$ be a matrix. We define $v(y)_{i_{0}}$ as the $i_{0}$-th column of $v(y)$. We define the function $v(y)_{i_{0}}$ maps a $d^{2}$-dimensional vector to an $n$-dimensional vector, given each $i_{0}$ in the set $[d]$, such that $v(y)_{i_{0}}:=A_{3}Y_{*,i_{0}}$ where $y\in\mathbb{R}^{d^{2}}$ is the vectorization of $n\times n$ matrix $Y$.
@@
Definition B.5 (RoPE attention error $\ell(x)$).
From Def. B.3, with every $j_{0}$ in the set $[n]$, it gives $s(x)_{j_{0}}$ as an $n$-dimensional normalized vector, and we define $v(y)_{i_{0}}$ based on Def. B.4 given that each $i_{0}\in[d]$. Defining a function $\ell(x)_{j_{0},i_{0}}$ maps a $d^{4}$-dimensional vector to a scalar with each $j_{0}\in[n]$ and each $i_{0}\in[d]$ such that
@@
Definition B.6 (Loss term $\mathsf{Loss}(x)$).
@@
Appendix C RoPE Attention Gradient Calculation
In this section, we analyze the time complexity of exact gradient computation. In Section C.1, we reformulate the closed form of the gradient. In Section C.2, we show the time complexity for $s(x)$ and $v(y)$. In Section C.3, we show the time complexity for $\ell(x)$. In Section C.4, we show the time complexity for $\beta(x)$ and $\gamma(x)$. In Section C.5, we show the total time complexity for computing the gradient of RoPE attention.